OpenStudy (anonymous):

If x represents the first two consecutive integers, express the sum of the two integers in terms of x

OpenStudy (anonymous):

by ur problem it seems x=n and x=n+1 so n+n+1=x+x=2x is it so??

OpenStudy (anonymous):

It should be 2x+2

OpenStudy (anonymous):

what do u mean by x represents the first two consecutive integers?

OpenStudy (anonymous):

im really not sure. thats just what my book says. im having a hard time understanding this .

OpenStudy (anonymous):

me too..the statement is not clear.

OpenStudy (anonymous):

how one variable can represent two integers simultaneously??

OpenStudy (anonymous):

im kinda lost. thanks anyhow

OpenStudy (anonymous):

you most welcome...

OpenStudy (anonymous):

and thanks for posting such an extra ordinary problem

OpenStudy (anonymous):

hello @myininaya what do u think of it??

myininaya (myininaya):

i think it means let x present the first integer of two consecutive intgers did you miss this word of?

OpenStudy (anonymous):

then if x=n then n+1=x+1 s0 sum is 2x+1 not 2x+2!!

myininaya (myininaya):

this would add meaning to the problem otherwise it makes no sense at all

OpenStudy (anonymous):

yes...

myininaya (myininaya):

go prepositions!

OpenStudy (anonymous):

@myininaya do u have any really tough questions regarding trigonometry or algebra..

myininaya (myininaya):

why you want me to give you one?

OpenStudy (anonymous):

i want try hard problems...if u have.

OpenStudy (anonymous):

then atleast we could disscuss it..

myininaya (myininaya):

let me think okay i will make a new post when i come up with one

OpenStudy (anonymous):

ok